OpenStudy (anonymous):

Solve the radical equation √(x+4) +√(x-4)=4

OpenStudy (anonymous):

I don't know where to begin...can I square everything? to get rid of the square roots, that is.

OpenStudy (anonymous):

√(x+4)=4 -√(x-4) x+4=x-8√(x-4)+12 4=-8√(x-4)+12

OpenStudy (anonymous):

think you can solve it from there

OpenStudy (anonymous):

I am confused....why did you do what you just did?

OpenStudy (anonymous):

subtract √(x-4) from both sides

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

I understand. but why did you have to subtract the √(x-4) from both sides? can't we just square everything and combine like terms, isolate x and that is our answer?

OpenStudy (anonymous):

well, if you square both sides, you get 2x+√(x+4)√(x-4)=16, which looks a lot more complicated

OpenStudy (anonymous):

I mean 2x+2√(x+4)√(x-4)=16

OpenStudy (anonymous):

no...you would get: (√(x+4))^2 +(√(x-4))^2=(4)^2 x+4 +x-4=16 which is 2x=16 x=8

OpenStudy (anonymous):

that is my logic anyway

OpenStudy (anonymous):

no, you gotta square everything

OpenStudy (anonymous):

but I did

OpenStudy (anonymous):

(√(x+4) +√(x-4))^2=(4)^2

OpenStudy (anonymous):

the square gets rid of the square rooots

OpenStudy (anonymous):

(√(x+4) +√(x-4))(√(x+4) +√(x-4))=16

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

(√(x+4) +√(x-4))(√(x+4) +√(x-4)) does not equal (√(x+4))^2 +(√(x-4))^2

OpenStudy (anonymous):

that is irritating... -_- so either way, I am going to have to foil....?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

meh :( well, thanks for your help

OpenStudy (anonymous):

np